# Deep Learning Accelerated Quantum Transport Simulations in Nanoelectronics: From Break Junctions to Field-Effect Transistors

### v1

@@ Abstract
Quantum transport calculations are essential for understanding and designing nanoelectronic devices, yet the trade-off between accuracy and computational efficiency has long limited their practical applications. We present a general framework that combines the deep learning tight-binding Hamiltonian (DeePTB) approach with the non-equilibrium Green’s Function (NEGF) method, enabling efficient quantum transport calculations while maintaining first-principles accuracy. We demonstrate the capabilities of the DeePTB-NEGF framework through two representative applications: comprehensive simulation of break junction systems, where conductance histograms show good agreement with experimental measurements in both metallic contact and single-molecule junction cases; and simulation of carbon nanotube field effect transistors through self-consistent NEGF-Poisson calculations, capturing essential physics including the electrostatic potential and transfer characteristic curves under finite bias conditions. This framework bridges the gap between first-principles accuracy and computational efficiency, providing a powerful tool for high-throughput quantum transport simulations across different scales in nanoelectronics.
@@ I Introduction
Quantum transport simulations are essential for understanding and designing nanoelectronic devices. The non-equilibrium Green’s function (NEGF) method[1, 2, 3], combined with density functional theory (DFT)[4, 5], has established itself as the standard approach for investigating quantum transport phenomena at the first-principles level[6, 7]. However, the computational burden of DFT-NEGF self-consistent field (SCF) iterations severely constrains its practical applications. This limitation becomes particularly critical in studying large-scale systems and dynamic processes, in which extensive sampling of configuration space is required. A prime example is the break junction experiments[8, 9], where quantum transport through nanoscale contacts evolves from metallic contacts to single-molecule junctions. Thousands of conductance measurements throughout this process are essential for statistical analysis, making first-principles calculations computationally prohibitive.
@@
Machine learning (ML) approaches have emerged as promising alternatives to accelerate quantum transport calculations. There are several ML attempts including direct prediction of conductance from atomic structures using local descriptors[10], and mapping combined structural and electronic features to transmission spectra near the Fermi level[11]. While these approaches demonstrate the potential of ML in transport calculations, they are often limited to specific systems or single transport property, lacking the generality required for broad applications in nanoelectronics. A systematic and comprehensive framework to accelerate DFT-NEGF simulations and capable of predicting various transport properties - from transmission spectra to potential profiles and current-voltage characteristics - remains elusive.
@@
The key to addressing this challenge lies in bypassing the time-consuming SCF iterations while maintaining first-principles accuracy. Our recently developed deep learning represented tight-binding (TB) Hamiltonian method, DeePTB[12, 13], provides an efficient route to accurate electronic structure calculations. By combining deep neural networks with physics-based principles, DeePTB achieves remarkable accuracy in predicting both local environment-dependent Slater-Koster (SK) tight-binding Hamiltonians[12] and DFT Hamiltonians, overlap and density matrix in the linear combination of atomic orbitals (LCAO) basis[13]. Building on these capabilities, we present a general DeePTB-NEGF framework for efficient quantum transport calculations. This framework enables direct prediction of TB models or DFT Hamiltonians as well as the overlap matrices, eliminating the need for SCF iterations while maintaining first-principles accuracy.
@@
We demonstrate the versatility and accuracy of our approach through two representative applications: (1) break junction systems, where we successfully simulate the dynamic process in metallic contacts and single-molecule junctions, showing good agreement with both DFT-NEGF calculations and experimental measurements[14]; and (2) carbon nanotube field-effect transistors (CNT-FETs), where we implement Poisson calculations in DeePTB-NEGF framework to study transport characteristics under finite bias, validating our results against the established NanoTCAD ViDES solver[15, 16]. These applications demonstrate that our framework can efficiently handle quantum transport across different scales and materials systems, opening new possibilities for high-throughput quantum transport calculations in nanoelectronics.
@@ II Methodology and Workflow
The key to efficient quantum transport calculations lies in obtaining accurate electronic structure information while avoiding expensive SCF iterations. Our approach combines the NEGF formalism for transport calculations with the DeePTB method for efficient Hamiltonian prediction, as illustrated in Fig. 1. DeePTB offers two complementary strategies for constructing electronic Hamiltonians, ie. the environment-dependent SK tight-binding Hamiltonians[12] and DFT Hamiltonians in the LCAO basis with E(3) equivariant graph neural networks[13].
@@
Figure 1: Schematic illustration of the DeePTB-NEGF workflow. Atomic configurations $X(t)$ are generated from molecular dynamics simulations. Reference Hamiltonians $\hat{\bm{H}}(X)$ and overlap matrices $\hat{\bm{S}}(X)$ are obtained from DFT-NEGF calculations, while $\bm{H}(X)$ and $\bm{S}(X)$ represent predictions from the DeePTB model. $T(E)$ denotes the transmission coefficient at energy $E$ and $I$ the current through the two-terminal device.
@@
where $f_{L/R}(E)$ are Fermi-Dirac distributions for the left and right electrodes, $E_{f}$ is the Fermi energy, and $\mathrm{G}_{0}=2e^{2}/h$ is the quantum conductance, $e$ is the elementary charge, and $h$ is the Planck’s constant.
@@
The complete workflow is illustrated in Fig. 1. Molecular dynamics (MD) simulations are used to generate atomic configurations $X(t)$ at finite temperatures along the trajectories at different time $t$, which are then divided into training and prediction sets. For the training set, configurations are labeled with converged Hamiltonians $\hat{\bm{H}}(X)$ and overlap matrices $\hat{\bm{S}}(X)$ from DFT-NEGF calculations using TranSIESTA[21, 6, 17]. The trained DeePTB model then efficiently predicts Hamiltonians $\bm{H}(X)$ and overlap matrices $\bm{S}(X)$ for new configurations, enabling high-throughput quantum transport calculations through the optimized NEGF framework.
@@ III.1 Break Junction Systems
Break junction experiments, primarily implemented through mechanically controllable break junctions (MCBJ) and scanning tunneling microscopy break junctions (STM-BJ)[8, 9], provide a powerful platform for investigating quantum transport at the atomic and molecular scales. During a typical break junction process, the configuration evolves from the metallic contacts to the molecular junctions. In the metallic contacts, conductance exhibits characteristic quantization features reflecting the discrete nature of atomic contacts. As the junction breaks, individual molecules can bridge the nanoscale gap forming a stable molecular junction[22]. Statistical analysis of these transport features requires thousands of conductance measurements over repeated breaking events, making first-principles calculations computationally prohibitive. The DeePTB-NEGF framework developed here enables efficient simulation of the break junction process while maintaining the first-principles accuracy. Below we demonstrate its capabilities in the two cases of metallic contacts and molecular junctions, providing direct comparison with experimental measurements and atomic-level insights into the quantum transport mechanisms.
@@ III.1.1 Quantum Transport in Metallic Contacts
To simulate quantum transport through gold contacts, we construct a junction model consisting of a central scattering region connected to two semi-infinite electrodes, as shown in the top panel of Fig. 2(a). The scattering region contains 304 gold atoms arranged along the [100] direction, with two bulk-like sections (100 atoms each in $5\times 5\times 4$ supercells) serving as electrode extensions. To simulate the break junction formation, we performed molecular dynamics simulations using the DeePMD potential[25] at 150K in the NVT ensemble, with the simulation box being the scattering region under periodic boundary conditions. During elongation, the electrode sections maintain their bulk structure while moving in opposite directions, with stretching speeds ranging from 0.2 to 5.0 m/s. The structural evolution exhibits characteristic bipyramidal shapes consistent with experimental high-resolution transmission electron microscopy observations of gold contacts under stretching [26].
For constructing the DeePTB model for gold contacts, we randomly sampled 122 configurations from 4 independent trajectories at $v=5$ m/s and calculated their electronic Hamiltonians and overlaps using DFT-NEGF with PBE functional and SZP basis in TranSIESTA. The model’s cutoff radius for the local environment of gold atom $r_{cut}^{\text{Au}}$ was set to 7.4Å to match the maximum radius of gold atomic orbitals in DFT calculations. Although the training configurations only include the scattering region, periodic boundary conditions ensure that atoms near the simulation box edges experience a bulk-like environment within the cutoff radius as shown in the lower panel of Fig. 2(a), crucial for accurate self-energy calculations in the NEGF framework. Details of data preparation and model training are provided in the supplementary materials (SM).
@@
We then validate the model’s predictive capability for electronic properties. Despite being trained only on scattering regions, the model successfully predicts Hamiltonian elements for larger systems shown in the top panel of Fig. 2(a) with additional electrode layers, exhibiting excellent agreement with DFT results ($\text{RMSE}=4.3\times 10^{-4}$ eV for Hamiltonian elements, see SM for details). More importantly, we examine the model’s generalization ability across different elongation speeds. Although trained solely on configurations from trajectories with $v=5.0$ m/s, the model demonstrates excellent transferability. By analyzing 55 randomly selected configurations spanning speeds from 0.2 m/s to 5.0 m/s, we compare zero-bias conductances predicted by DeePTB-NEGF with DFT-NEGF results. The model achieves remarkable accuracy with $R^{2}=0.9985$ in the parity plot and maintains low prediction errors (RMSE and MAE $<$ $60\times 10^{-3}\mathrm{G}_{0}$) across all elongation speeds, as shown in Fig. 2(c) and (d).
@@
Having validated the model’s accuracy, we then applied the model to calculate the gold contact conductance and demonstrate its efficiency in high-throughput calculations. Using the DeePTB model trained only on 122 configurations, we successfully generated conductance histograms by analyzing 9224 configurations sampled from 105 elongation processes at speeds of 0.2, 1.0, and 5.0 m/s. The calculations were accelerated using the Bloch theorem for self-energy evaluation and exploiting the tridiagonal block structure of Hamiltonian matrices. The histograms, normalized to their respective $1\mathrm{G}_{0}$ peaks, are presented in Fig.2(e). The prominent peak at $1\mathrm{G}_{0}$ corresponds to monatomic chains, a well-established signature in metallic break junctions.
@@ III.1.2 Quantum Transport through Single-molecule Junctions
The molecule investigated here is a classical $\pi$-conjugated system with sulfur methyl as anchoring groups, which contains rich chemical environments and flexible configurations due to three types of carbon bonds and the dihedral angle between the two benzene planes (top left inset of Fig. 3(c)). The scattering region consists of the molecule connected to gold electrodes with a bi-pyramidal shape, where each electrode principal layer contains 144 atoms ($6\times 6\times 4$ supercell). To simulate the breaking junction process, we performed MD simulations at 300K using the ReaxFF reactive force field[28], which can efficiently describe chemical bonding through its empirical bond-order formalism. During stretching, the gold electrodes move in opposite directions at a speed of $v=\pm 2$ m/s, with three representative snapshots at $t_{1}$, $t_{2}$, $t_{3}$ in equal time intervals shown in Fig. 3(a). To construct the DeePTB model for molecular junctions, we randomly sampled 268 configurations from 14 breaking junction trajectories as the training set and calculated their electronic Hamiltonians and overlaps using DFT-NEGF with PBE functional and SZP basis in TranSIESTA. Similarly, the model’s cutoff radius was set as the largest radius of atomic orbitals for each element in DFT-NEGF calculations, ensuring accurate description of electronic interactions. More details of data preparation and model training are provided in the SM.
@@
III.2 Double-gated Carbon Nanotube Field-effect Transistor
Figure 4: DeePTB-NEGF-Poisson SCF simulation results for double-gated carbon nanotube field-effect transistor. (a) Schematic of the transistor geometry, where $L_{\text{G}}$ denotes gate length and $t_{\text{ox}}$ represents gate dielectric thickness. (b) Transmission spectrum at gate voltage $V_{\text{gs}}=1$ V and drain-source bias $V_{\text{ds}}=0.5$ V. (c) Electrostatic potential energy distribution in the $xz$ plane. (d) Difference in electrostatic potential energy between DeePTB-NEGF and NanoTCAD ViDES calculations in the $xz$ plane, using identical device geometry and boundary conditions ($L_{\text{G}}=2$ nm, $V_{\text{gs}}=1$ V). (e) Transfer characteristics comparing DeePTB-NEGF and NanoTCAD ViDES results for SiO2 and HfO2 dielectrics with $L_{\text{G}}=2$ nm and $V_{\text{ds}}=0.5$ V. (f) Transfer characteristics comparing DeePTB-NEGF and NanoTCAD ViDES results for different gate lengths ($L_{\text{G}}=2$ nm and $10.5$ nm) with SiO2 dielectric and $V_{\text{ds}}=0.5$ V.

### v2

@@ Abstract
Quantum transport simulations are essential for understanding and designing nanoelectronic devices, yet the longstanding trade-off between accuracy and computational efficiency has limited their practical applications. We present a general integrated framework that combines the deep learning tight-binding Hamiltonian (DeePTB) approach with the non-equilibrium Green’s Function (NEGF) method, enabling efficient quantum transport calculations with first-principles accuracy. The capabilities of the DeePTB-NEGF framework are demonstrated through two representative applications: comprehensive simulation of break junction systems, where conductance histograms exhibit good agreement with experimental measurements in both metallic contact and single-molecule junction cases; and simulation of carbon nanotube field-effect transistors in the experimental dimensions using NEGF-Poisson self-consistency, achieving excellent alignment with measured transfer characteristics. Furthermore, our simulation reveals the limited scaling down potential for local bottom gate geometry into sub-5 nm due to performance degradation. By bridging the gap between first-principles accuracy and computational efficiency, our framework provides a powerful tool for high-throughput quantum transport simulations across different scales in nanoelectronics.
@@ I Introduction
Quantum transport simulations are essential for understanding and designing nanoelectronic devices. The non-equilibrium Green’s function (NEGF) method[1, 2, 3], combined with density functional theory (DFT)[4, 5], has established itself as the standard approach for investigating quantum transport phenomena at the first-principles level[6, 7]. However, the computational burden of DFT-NEGF self-consistent field (SCF) iterations severely constrains its practical applications, which is particularly critical in studying large-scale systems and dynamic processes. This limitation is especially critical in scenarios such as break junction experiments [8, 9], where thousands of conductance measurements are required for statistical analysis, and in simulating field-effect transistors (FETs) with experimentally realistic dimensions. In both cases, the extensive computational resources required make first-principles calculations prohibitively expensive, often restricting studies to selected configurations or significantly downsized systems. This compromise between computational feasibility and experimental relevance highlights the urgent need for more efficient approaches that can maintain first-principles accuracy while dramatically reducing computational costs.
@@
Machine learning (ML) approaches have emerged as promising alternatives to accelerate quantum transport calculations. There are several ML attempts including direct prediction of conductance from atomic structures using local descriptors[10], and mapping combined structural and electronic features to transmission spectra near the Fermi level[11]. While these approaches demonstrate the potential of ML in transport calculations, they are often limited to specific systems or single transport property, lacking the generality required for broad applications in nanoelectronics. A systematic and comprehensive framework to accelerate DFT-NEGF simulations and capable of predicting various transport properties - from transmission spectra to current-voltage characteristics - remains elusive.
@@
The key to addressing this challenge is to bypass the time-consuming DFT-NEGF SCF iterations while maintaining first-principles accuracy. Our recently developed deep learning represented tight-binding (TB) Hamiltonian method, DeePTB[12, 13], provides an efficient route to accurate electronic structure calculations. By combining deep neural networks with physics-based principles, DeePTB achieves remarkable accuracy in predicting both local environment-dependent Slater-Koster (SK) tight-binding Hamiltonians[12] and DFT Hamiltonians, overlap and density matrix in the linear combination of atomic orbitals (LCAO) basis[13]. Building on these capabilities, we introduce DeePTB-NEGF, a general integrated framework for efficient quantum transport calculations. This framework enables direct prediction of TB models or DFT Hamiltonians as well as the overlap matrices, eliminating the need for DFT-NEGF SCF while preserving first-principles accuracy.
@@
We demonstrate the versatility and accuracy of our approach through two representative applications: (1) break junction systems, where we successfully simulate the dynamic process of metallic contacts and single-molecule junctions, showing good agreement with both DFT-NEGF calculations and experimental measurements[14]; and (2) carbon nanotube field-effect transistors (CNT-FETs), where we incorporate Poisson equation calculations within the DeePTB-NEGF framework to model the transport characteristics of CNT-FETs at experimental-scale dimensions under finite bias, exhibiting excellent consistency with the experimental results [15]. These applications demonstrate that our framework can efficiently handle quantum transport across different scales and materials systems, opening new possibilities for high-throughput quantum transport calculations in nanoelectronics.
@@ II Methodology and Workflow
Efficient first-principles quantum transport simulations require accurate electronic structure description while avoiding the computationally expensive DFT-NEGF SCF iterations. Our approach combines the NEGF formalism for transport calculations with the DeePTB method for efficient Hamiltonian prediction, as illustrated in Fig. 1. DeePTB offers two complementary strategies for constructing electronic Hamiltonians: (1) the environment-dependent SK tight-binding Hamiltonians[12] and (2) DFT Hamiltonians in the LCAO basis with E(3) equivariant graph neural networks[13].
@@
Figure 1: Schematic illustration of DeePTB-NEGF workflow. Atomic configurations $X$ and $X^{\prime}$ are structures for training and prediction respectively. Reference Hamiltonians/overlap matrices $\hat{\bm{H}}(X)$/$\hat{\bm{S}}(X)$ and reference eigenvalues $\hat{\epsilon}_{\text{nk}}(X)$ are obtained from first-principles calculations, while $\bm{H}(X^{\prime})$/ $\bm{S}(X^{\prime})$ and $\bm{H}_{\text{TB}}(X^{\prime})$ represent predictions from DeePTB model with E3 method and SK method respectively. $T(E)$ denotes the transmission coefficient at energy $E$ and $I$ the current through the device.
@@
where $f_{L/R}(E)$ are Fermi-Dirac distributions for the left and right electrodes, $E_{f}$ is the Fermi energy, $\mathrm{G}_{0}=2e^{2}/h$ is the quantum conductance, $e$ is the elementary charge, and $h$ is the Planck’s constant.
@@
The complete workflow is illustrated in Fig. 1. The training set consists of atomic configurations labeled with converged Hamiltonians and overlap matrices $\hat{\bm{H}}(X)$ and $\hat{\bm{S}}(X)$ obtained from DFT-NEGF calculations using TranSIESTA[6, 16], or eigenvalues $\hat{\epsilon}_{\text{nk}}(X)$ from DFT calculations performed with SIESTA[20].For a new configuration $X^{\prime}$, the trained DeePTB model can efficiently predict the DFT Hamiltonians and overlap matrices $\bm{H}(X^{\prime})$ and $\bm{S}(X^{\prime})$, or construct tight binding Hamiltonians $\bm{H}_{\text{TB}}(X^{\prime})$, enabling high-throughput quantum transport calculations through the optimized NEGF framework. Furthermore, the DeePTB-NEGF framework extends to quantum transport simulations in devices under finite bias via NEGF-Poisson SCF iterations, as detailed in Section III.2.
@@ III.1 Break Junction Systems
Break junction experiments, primarily implemented through mechanically controllable break junctions or scanning tunneling microscopy break junctions[8, 9], provide a powerful platform for investigating quantum transport at the atomic and molecular scales. During a typical break junction process, the configuration evolves from the metallic contacts to the molecular junctions. In the metallic contacts, conductance exhibits characteristic quantization features reflecting the discrete nature of atomic contacts. As the junction breaks, individual molecules can bridge the nanoscale gap forming a stable molecular junction[21]. Statistical analysis of these transport features requires thousands of conductance measurements over repeated breaking events, making first-principles calculations computationally prohibitive. The DeePTB-NEGF framework developed here enables efficient simulation of the break junction process while maintaining the first-principles accuracy. Below we demonstrate its capabilities in the two cases of metallic contacts and molecular junctions, providing direct comparison with experimental measurements and atomic-level insights into the quantum transport mechanisms.
@@ III.1.1 Quantum Transport in Metallic Contacts
To simulate quantum transport through gold contacts, we construct a junction model consisting of a central scattering region connected to two semi-infinite electrodes, as shown in the top panel of Fig. 2(a). The scattering region contains 304 gold atoms arranged along the [100] direction, with two bulk-like sections (100 atoms each in $5\times 5\times 4$ supercells) serving as electrode extensions. To simulate the break junction formation, we performed molecular dynamics simulations using the DeePMD potential[24] at 150K in the NVT ensemble, with the simulation box being the scattering region under periodic boundary conditions. During elongation, the electrode sections maintain their bulk structure while moving in opposite directions, with stretching speeds ranging from 0.2 to 5.0 m/s in each side. The structural evolution exhibits characteristic bipyramidal shapes consistent with experimental observations of gold contacts by high-resolution transmission electron microscopy [25].
For constructing the DeePTB model for gold contacts, we randomly sampled 122 configurations from 4 independent trajectories at $v=\pm 5$ m/s and calculated their electronic Hamiltonians and overlaps using DFT-NEGF with generalized gradient approximation (GGA) within the Perdew-Burke-Ernzerhof (PBE) formulation as exchange-correlation functional and single-zeta plus polarization(SZP) basis. The model’s cutoff radius for the local environment of gold atom $r_{cut}^{\text{Au}}$ was set to 7.4Å to match the maximum radius of gold atomic orbitals in DFT-NEGF calculations. Although the training configurations only include the scattering region, periodic boundary conditions utilized in DeePTB ensure that atoms near the simulation box edges experience a bulk-like environment within the cutoff radius, crucial for accurate self-energy calculations in the NEGF framework. As shown in the lower panel of Fig. 2(a) and supplementary materials (SM) Fig. S1, in regions at simulation box edge (far from the structure interface), the electronic Hamiltonian matches that of bulk gold with identical local environments, differing only by a constant shift due to long-range effects. We validated this phenomenon by comparing DeePTB-predicted and DFT-calculated Hamiltonians and overlap matrices for edge atom with its bulk counterpart. The results confirm the expected shift and demonstrate the model’s accuracy for electrodes. More details of data preparation, model training and validation of model in bulk-like environment are available in SM Sec.S1.A.
@@
We then validate the model’s predictive capability for electronic properties in complete gold atomic contacts, as shown in the top panel of Fig. 2(a). Building on its accuracy in bulk-like environment, the model successfully predicts Hamiltonian for the complete atomic contact including additional electrode layers, exhibiting excellent agreement with DFT results (Hamiltonian $\text{RMSE}=7.57\times 10^{-4}$ eV for configurations in validation set, see SM Sec.S1.A for details). More importantly, we assess the model’s generalization ability across different elongation speeds. Although trained exclusively on configurations from trajectories with $v=5.0$ m/s, the model demonstrates outstanding transferability. By analyzing 55 randomly selected configurations spanning speeds from 0.2 m/s to 5.0 m/s, we compare zero-bias conductances predicted by DeePTB-NEGF with DFT-NEGF results. The model achieves exceptional accuracy with $R^{2}=0.9985$ in the parity plot and maintains low prediction errors (conductance RMSE and MAE $<$ $60\times 10^{-3}\mathrm{G}_{0}$) across all elongation speeds, as shown in Fig. 2(c) and (d).
@@
Having validated the model’s accuracy, we then applied the model to calculate the gold contact conductance and demonstrate its efficiency in high-throughput calculations. Using the DeePTB model trained only on 122 configurations randomly sampled from 4 breaking junction trajectories, we successfully generated conductance histograms by analyzing 9224 configurations randomly sampled from 105 elongation processes at speeds of 0.2, 1.0, and 5.0 m/s. The calculations were accelerated using the Bloch theorem for self-energy evaluation and exploiting the tridiagonal block structure of Hamiltonian matrices. The histograms, normalized to their respective $1\mathrm{G}_{0}$ peaks, are presented in Fig.2(e). The prominent peak at $1\mathrm{G}_{0}$ corresponds to monatomic chains, a well-established signature in metallic break junctions.
@@ III.1.2 Quantum Transport through Single-molecule Junctions
The molecule investigated here is a classical $\pi$-conjugated system with sulfur methyl as anchoring groups, which contains rich chemical environments and flexible configurations due to three types of carbon bonds and the dihedral angle between the two benzene planes (top left inset of Fig. 3(c)). The scattering region consists of the molecule connected to gold electrodes with a bi-pyramidal shape, where each electrode principal layer contains 144 atoms ($6\times 6\times 4$ supercell). To simulate the breaking junction process, we performed MD simulations at 300K using the ReaxFF reactive force field[27], which can efficiently describe chemical bonding through its empirical bond-order formalism. During stretching, the gold electrodes move in opposite directions at a speed of $v=\pm 2$ m/s, with three representative snapshots at $t_{1}$, $t_{2}$, $t_{3}$ in equal time intervals shown in Fig. 3(a). To construct the DeePTB model for molecular junctions, we randomly sampled 268 configurations from 14 breaking junction trajectories as the training set and calculated their electronic Hamiltonians and overlaps using DFT-NEGF with PBE functional and SZP basis in TranSIESTA. Similarly, the model’s cutoff radius was set as the largest radius of atomic orbitals for each element in DFT-NEGF calculations, ensuring accurate description of electronic interactions. More details of data preparation and model training are provided in the SM Sec.S1.B.
@@
III.2 Carbon Nanotube Field-effect Transistor
Figure 4: DeePTB-NEGF-Poisson SCF simulation for local bottom gate carbon nanotube (LBG) field-effect transistor. (a) Schematic of the transistor geometry, where $L_{\text{ch}}$ denotes channel length. (b)Transfer characteristics at drain-source bias $V_{\text{ds}}=0.4$ V and dope concentrations as $8.89\times 10^{8}\text{m}^{-1}$ for LBG CNT FETs using different CNT diameters. The channel length $L_{\text{ch}}$ for all FETs are 9nm.(c)Transfer characteristics at drain-source bias $V_{\text{ds}}=0.4$ and dope concentrations as $8.89\times 10^{8}\text{m}^{-1}$ for LBG CNT FETs using different channel lengths. The tubes are all CNT$(16,0)$. (d) Position-resolved local density of states for CNT$(17,0)$ along the transport direction($z$ axis) with gate voltage $V_{\text{gs}}=-0.8\text{V},-0.4\text{V},0.0\text{V}$. The left and right electrodes are drain and source respectively. The red dashed lines indicate the Fermi level for source and drain. The orange dashed line in $V_{\text{gs}}=0.0$ case indicates the top of valence band inducing a barrier $\Phi_{B}=0.19$ eV for holes.